# Temporal Action Localization for Inertial-based Human Activity Recognition

### v1

@@ Abstract.
A persistent trend in Deep Learning has been the applicability of machine learning concepts to other areas than originally introduced for. As of today, state-of-the-art activity recognition from wearable sensors relies on classifiers being trained on fixed windows of data. Contrarily, video-based Human Activity Recognition has followed a segment-based prediction approach, localizing activity occurrences from start to end. This paper is the first to systematically demonstrate the applicability of state-of-the-art TAL models for wearable Human Activity Recongition (HAR) using raw inertial data as input. Our results show that state-of-the-art TAL models are able to outperform popular inertial models on 4 out of 6 wearable activity recognition benchmark datasets, with improvements ranging as much as $25\%$ in F1-score. Introducing the TAL community’s most popular metric to inertial-based HAR, namely mean Average Precision, our analysis shows that TAL models are able to produce more coherent segments along with an overall higher NULL-class accuracy across all datasets. Being the first to provide such an analysis, the TAL community offers an interesting new perspective to inertial-based HAR with yet to be explored design choices and training concepts, which could be of significant value for the inertial-based HAR community.
@@ 1. Introduction
The recognition of performed activities through wearable sensors such as IMUs has shown to be of significant value in areas such as health care or the support of complex work processes (Bulling et al., 2014). Deep Learning has established itself as the de facto standard in wearable, inertial-based Human Activity Recognition (HAR), consistently outperforming classical Machine Learning approaches in recognition performance. A persistent trend in Deep Learning has been the applicability of machine learning concepts such as self-attention (Vaswani et al., 2017) to other areas and application scenarios than originally introduced for. With significant progress having been made since the introduction of deep neural architectures such as the DeepConvLSTM (Ordóñez and Roggen, 2016), researchers have followed this trend and continuously worked on improving the architectural design of networks by incorporating newly introduced techniques (see e.g., (Zhou et al., 2022)).
A promising recent approach in video-based Human Activity Recognition (HAR) is Temporal Action Localization (TAL), which aims to locate activity segments, defined by a class label, start, and end point, within an untrimmed video. Even though introduced architectures have almost doubled in performance over the last 5 years on existing datasets like THUMOS-14 (Jiang et al., 2014), results on large corpora such as EPIC-KITCHENS-100 (Damen et al., 2022) and Ego4D (Grauman et al., 2022) show that the prediction problem is far from being saturated. Recently, Zhang et al. (2022b) introduced the ActionFormer, an end-to-end, single-stage TAL model utilizing transformer-based layers. Being the first of its kind, the ActionFormer along with extensions of it (Tang et al., 2023; Shi et al., 2023) are as of today holding top ranks in prediction performance on a multitude of TAL benchmark datasets. Despite sharing a mutual goal, the TAL community evaluates their approaches differently from the inertial community, using mean Average Precision (mAP) as opposed to traditional classification metrics like F1-score or accuracy. TAL models have recently been shown to be capable of being trained using raw inertial data (Bock et al., 2023), marking the first instance of such vision-based models being applied in the context of inertial-based HAR. When compared with popular inertial models, results on one particular dataset have already shown that TAL models can produce more coherent and less fragmented predictions while maintaining performance in terms of traditional classification metrics.
Setting out to further examine the applicability of Temporal Action Localization for inertial-based HAR, our contributions in this paper are three-fold:
@@
    We demonstrate the capabilities of a novel approach to inertial-based HAR, being single-stage TAL, to outperform popular inertial models on four out of six wearable activity recognition benchmark datasets.
@@
    Complementing traditional, frame-wise classification metrics, we introduce a set of unexplored, segment-based evaluation metrics for inertial-based HAR, which are based around the scalar evaluation metric mean Average Precision (Alwassel et al., 2018).
@@
    We name differences and similarities between models originating from both communities, highlighting applied concepts that TAL models use to produce coherent segments along with a high NULL-class accuracy.
@@ 2. Related Work
With on-body sensors providing a robust and non-intrusive way to monitor participants along long stretches of time, research conducted in the area of inertial-based HAR has worked towards the automatic interpretation of one or multiple sensor streams to reliably detect activities e.g. in the context of providing medical support or providing guidance during complex work processes (Bulling et al., 2014). With deep neural networks having established themselves as the de facto standard in inertial-based HAR, the DeepConvLSTM (Ordóñez and Roggen, 2016) as well as models introduced at a later point all follow a similar prediction scenario design applying a sliding window approach which groups concurrent data points for classification. In their original publication Ordóñez and Roggen (2016) have found a combination of convolutional and recurrent layers to produce competitive results, with the idea being to model temporal dependencies amongst automatically extracted discriminative features within a sliding window in order to classify it correctly as one of the $N$ activity classes or, if applicable, NULL-class. Building up on the idea of combining these two types of layers, researchers have worked on extending the original DeepConvLSTM or have introduced their own architectural designs (Murahari and Plötz, 2018; Vaizman et al., 2018; Peng et al., 2018; Xi et al., 2018; Xu et al., 2019; Xia et al., 2020; Liu et al., 2020; Chen et al., 2020; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022; Zhang et al., 2022a; Miao et al., 2022; Chen et al., 2023b, a; Wieland and Pankratius, 2023). A simple modification of the DeepConvLSTM is the shallow DeepConvLSTM (Bock et al., 2021). Contradicting the popular belief that one needs at least two recurrent layers when dealing with time series data (Karpathy et al., 2015), Bock et al. (2021) demonstrated that removing the second LSTM layer within the original DeepConvLSTM architecture, results in significant increases in predictive performance on a multitude of HAR benchmark datasets while also decreasing the number of learnable parameters. Furthermore, with the original DeepConvLSTM only being able to learn temporal dependencies within a sliding window, the shallow DeepConvLSTM applies the LSTM across batches, effectively making the batches the sequence which is to be learned by the LSTM. This dimension-flip, along with a non-shuffled training dataset, enables the architecture to learn temporal dependencies amongst a batched input. The same year as the shallow DeepConvLSTM, Abedin et al. (2021) introduced Attend-and-Discriminate, a deep neural network architecture following the idea of the original DeepConvLSTM by combining both convolutional and recurrent layers, yet further adding a cross-channel interaction encoder using self-attention as well as attention mechanism to the recurrent parts of the network. In 2022 Zhou et al. (2022) proposed TinyHAR, a lightweight HAR model that uses a transformer encoder block along with means of cross-channel fusion via a fully connected layer and a final self-attention layer which aims to learn the contribution of each outputted time step produced by the recurrent parts individually.
@@
Classifying videos in the context of Human Activity Recognition can broadly be categorized into three main application scenarios: Action Recognition, which aims to classify trimmed videos into one of $C$ activity classes; Action Anticipation, which aims to predict the next likely activities after observing a preceding video sequence; and Temporal Action Localization (TAL), which seeks to identify and locate all activity segments within an untrimmed video. With the inertial-based benchmark datasets consisting of a multitude of activities, TAL is to be considered most comparable to inertial-based HAR. Unlike popular intertial architectures though, TAL models aim to predict all segments within an untrimmed video at once. Existing TAL methods can broadly be categorized into two categories: two-stage and one-stage recognition systems. Two-stage recognition system (Lin et al., 2019, 2020; Xu et al., 2020; Bai et al., 2020; Zhao et al., 2020; Zeng et al., 2019; Gong et al., 2020; Liu et al., 2021; Qing et al., 2021; Sridhar et al., 2021; Zhu et al., 2021; Zhao et al., 2021; Tan et al., 2021) divide the process of identifying actions within an untrimmed video into two subtasks. First, during proposal generation, candidate segments are generated, which are then, during the second step, classified into one of $C$ activity classes and iteratively refined regarding their start and end times. Contrarily, single-stage models (Yang et al., 2023; Shi et al., 2022; Nag et al., 2022; Liu et al., 2022; Liu and Wang, 2020; Long et al., 2019; Lin et al., 2021; Chen et al., 2022; Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) do not rely on activity proposals, directly aiming to localize segments along their class label and refined boundaries. In 2022 Zhang et al. (2022b) introduced the ActionFormer, a lightweight, single-stage TAL model which unlike previously introduced single-stage architectures does not rely on pre-defined anchor windows. In line with the success of transformers in other research fields, Zhang et al. (2022b) demonstrated their applicability for TAL, outperforming previously held benchmarks on several TAL datasets (Heilbron et al., 2015; Damen et al., 2022; Jiang et al., 2014) by a significant margin. Surprisingly, a year later the TemporalMaxer suggested removing transformer layers within the ActionFormer arguing feature embeddings are already discriminative enough (Tang et al., 2023). Though being more lightweight than the ActionFormer, the TemporalMaxer showed to outperform its precedent during benchmark analysis. Similar to the TemporalMaxer, Shi et al. (2023) introduced TriDet, which suggested altering the ActionFormer in two ways. First, to mitigate the risk of rank-loss, self-attention layers are replaced with SGP layers. Second, the regression head in the decoder is replaced with a Trident head, which improves imprecise boundary predictions via an estimated relative probability distribution around each timestamp’s activity boundaries.
@@ 3. Temporal Action Localization for Inertial-based HAR
As the inertial-based HAR and TAL communities deal with inherently different modalities, both communities have developed distinct predictive pipelines and algorithms tailored to the challenges and characteristics of their respective modalities (see Figure 1). In the following section, we will provide detailed descriptions of the two different predictive pipelines and demonstrate how vision-based TAL models can be applied to a previously unexplored modality — inertial data. Furthermore, we provide a detailed overview of the architectural design of three popular single-stage TAL models (Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) highlighting similarities as well as differences of the architectures compared to models originating from the inertial-based HAR community.
3.1. Inertial-based HAR vs. Temporal Action Localization
@@
Problem definition
The objective of both inertial activity recognition and TAL is to predict all activities present in an untrimmed timeline. Given an input data stream $X$ of a sample participant, both the inertial and TAL communities start by applying a sliding window approach which shifts over $\mathbf{X}$, dividing the input data into windows, e.g. of one second duration with a 50% overlap between consecutive windows. This process results in $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$ being discretized into $t=\{0,1,...,T\}$ time steps, where $T$ is the number of windows for each participant. It is important to note that the TAL models do not use raw data as input, but are instead trained using feature embeddings extracted from each individual sliding video clip, which are extracted using pretrained methods such as (Carreira and Zisserman, 2017). That is, in the case of TAL, $\mathbf{x}_{t}$ represents a one-dimensional feature embedding vector associated with the video clip at time step $t$. In case of inertial-based activity recognition, $\mathbf{x}_{t}$ represents a two-dimensional matrix of size $\mathbb{R}^{W\times S}$, where $W$ is the number of samples within a window, and $S$ is the number of sensor axes present in the sensor data.
Given all sliding windows associated with an untrimmed sequence, inertial activity recognition models aim to predict an activity label $a_{t}$ for each sliding window $\mathbf{x}_{t}$, where $a$ belongs to a predefined set of activity labels, $a_{t}\in{1,...,C}$. To do so, the sliding windows are batched and fed through e.g. a deep neural network, such as the DeepConvLSTM (Ordóñez and Roggen, 2016). The resulting activity labels for each window are then compared to the true activity labels from the ground truth data, and classification metrics like accuracy or F1-score are calculated. Contrarily, TAL models aim to identify and localize segments of actions within the untrimmed data stream, which can span across multiple windows. To achieve this algorithms are trained to predict activity segments $Y=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$ using a collection of feature embeddings $X$ as input, where $N$ varies across participants. Each activity instance $\mathbf{y_{i}}=(s_{i},e_{i},a_{i})$ is defined by its starting time $s_{i}$ (onset), ending time $e_{i}$ (offset) and its associated activity label $a_{i}$, where $s_{i}\in[0,T]$, $e_{i}\in[0,T]$ and $a_{i}\in\{1,...,C\}$.
Evaluation metrics
As predictive outcomes of inertial activity recognition and TAL differ, both communities have adopted different evaluation metrics. In case of the inertial community, the problem of recognizing activities is translated to a per-window classification problem, where each window represents an instance to be classified into one of $C$ classes. Therefore, most researchers assess their algorithms based on traditional classification measures, including recall ($R$), precision ($P$), and F1-score ($F1$). In order to calculate these activity-wise metrics one first needs to calculate activity-wise True Positives (TP), False Positives (FP), True Negatives (TN) and False Negatives (FN). For a given activity class $i\in\{1,...,C\}$, a predicted activity label is to be treated as:
  • TP of class $i$ ($TP_{i}$): the predicted and ground truth label are equal to $i$
  • FP of class $i$ ($FP_{i}$): the predicted label is equal to $i$, yet the ground truth is not
  • TN of class $i$ ($TN_{i}$): the predicted and ground truth label are not equal to $i$
  • FN of class $i$ ($FN_{i}$): the predicted label is not equal to $i$, yet the ground truth is
Calculating TP, FP, TN and FN counts of all activities then allows to calculate the per-class classification metrics, which are defined as:
(1) $P_{i}=\frac{TP_{i}}{TP_{i}+FP_{i}}$
(2) $R_{i}=\frac{TP_{i}}{TP_{i}+FN_{i}}$
(3) $F1_{i}=\frac{2\cdot P_{i}\times R_{i}}{P_{i}+R_{i}}$
To account for the continuous nature of activity recognition prediction streams, Ward et al. (2011) introduced additional performance metrics to explicitly measure misalignment within predictions made by inertial-based activity recognition systems. According to the authors, calculating classification metrics such as precision and recall based on window-wise errors does not provide intuitive insights about the overall fragmentation and merging behavior of the prediction stream. The authors propose to further divide the class-wise counts of FPs and FNs into additional sub-categories based on their relative position within a segmented predictive stream. These misalignment sub-categories are:
  1. (1)
    Underfill ($U_{i}$): False Negatives of class $i$ that occur at the start or end of an activity segment, effectively making the predicted segment shorter by starting later or finishing earlier than the corresponding ground truth segment of $i$.
  2. (2)
    Deletion ($D_{i}$): False Negatives of class $i$ that correspond to a deleted event, i.e. a ground segment which is missed completely within the prediction stream.
  3. (3)
    Fragmenting ($F_{i}$): False Negatives of class $i$ that occur in-between True Positive instances of class $i$, falsely splitting a coherent segment into two segments.
  4. (4)
    Insertion ($I_{i}$): False Positives of class $i$ that correspond to a non-existent ground truth segment.
  5. (5)
    Overfill ($O_{i}$): False Positives of class $i$ that occur at the start or end of an activity segment, effectively making the predicted segment longer by predicting an earlier start or later finish than the corresponding ground truth segment.
  6. (6)
    Merge ($M_{i}$): False Positives of class $i$ that occur between two ground truth segments incorrectly merging the two segments into one coherent segment.
Calculating these six misalignment $U$, $D$, $F$, $I$, $O$ and $M$ counts of all activities then allows calculating the per-class misalignment ratios, which are defined as:
(4) $UR_{i}=\frac{U_{i}}{TN_{i}+FN_{i}}$
(5) $DR_{i}=\frac{D_{i}}{TN_{i}+FN_{i}}$
(6) $FR_{i}=\frac{F_{i}}{TN_{i}+FN_{i}}$
(7) $IR_{i}=\frac{I_{i}}{TP_{i}+FP_{i}}$
(8) $OR_{i}=\frac{O_{i}}{TP_{i}+FP_{i}}$
(9) $MR_{i}=\frac{M_{i}}{TP_{i}+FP_{i}}$
Inspired by measures applied during object detection, the TAL community assess algorithms based mean Average Precision (mAP) performance applied at different temporal Intersection over Union (tIoU) thresholds. With a TAL model being trained to predict a set of activity segments $\mathbf{Y}$, each predicted segment is compared with the ground truth activity segments within the complete input data stream of a participant. That is, given a predicted activity segment $\mathbf{y_{p}}=(s_{p},e_{p},a_{p})$ and a ground truth activity segment $\mathbf{y_{g}}=(s_{g},e_{g},a_{g})$, the two segments are considered to be either TP, FP, TN or FN based on the condition that the temporal intersection over union (tIoU) (overlap of the two segments divided by the union of the two segments) is greater than a predefined threshold, e.g., 0.5. Mean Average Precision is then calculated as:
(10) $mAP=\frac{1}{N}\sum^{C}_{a=1}{AP_{a}},$
where $AP_{i}$ represents the area under the curve of the Precision-Recall-Curve, which is obtained by plotting a model’s precision and recall values as a function of the model’s confidence score threshold for a given class $a\in\{1,...,C\}$. More specifically, the mAP score at a tIoU threshold of, for example, 0.2 (mAP@0.2), is calculated as the mAP score assuming a required tIoU of 0.2 when calculating $TP$, $FP$, $TN$ and $TN$ counts. Typically, to assess the predictive performance of TAL models, mAP is calculated using a range of tIoU thresholds, which are then averaged. In case of this benchmark analysis we calculate mAP as mAP@(0.3:0.1:07), which equates to calculating the mAP@0.3, mAP@0.4, mAP@0.5, mAP@0.6 and mAP@0.7 and averaging the individual scores.
3.2. Vectorization of Inertial Data for TAL
As previously mentioned, the input data $X$ used to train TAL models is a collection of feature embeddings. These embeddings are, for instance, extracted using some pre-trained method applied on each video clip. The resulting embeddings are 1-dimensional feature vectors that summarize information such as RGB and optical flow of the video clip (see, for example, Carreira and Zisserman (2017)). Since both communities employ a sliding window approach but feed data to their models using different dimensions, Bock et al. (2023) proposed a simple, yet effective preprocessing step. This step converts the 2-dimensional, windowed inertial data, as used by inertial architectures, into a 1-dimensional feature embedding suitable for training TAL models. The preprocessing method as detailed in Figure 2 involves concatenating the different sensor axes of each window, converting the input data to be a collection of 1-dimensional feature embedding vectors $\mathbf{x}_{t}\in\mathbb{R}^{1\times WS}$ where $W$ is the number of samples within a window, and $S$ is the number of sensor axes. More specifically, given a sliding window $\mathbf{x}_{SW}\in\mathbb{R}^{W\times S}$, we vectorize the two-dimensional matrix as follows:
@@ 3.2. Vectorization of Inertial Data for TAL
3.3. Architectures Overview

### v2

@@ Abstract.
As of today, state-of-the-art activity recognition from wearable sensors relies on algorithms being trained to classify fixed windows of data. In contrast, video-based Human Activity Recognition, known as Temporal Action Localization (TAL), has followed a segment-based prediction approach, localizing activity segments in a timeline of arbitrary length. This paper is the first to systematically demonstrate the applicability of state-of-the-art TAL models for both offline and near-online Human Activity Recognition (HAR) using raw inertial data as well as pre-extracted latent features as input. Offline prediction results show that TAL models are able to outperform popular inertial models on a multitude of HAR benchmark datasets, with improvements reaching as much as 26% in F1-score. We show that by analyzing timelines as a whole, TAL models can produce more coherent segments and achieve higher NULL-class accuracy across all datasets. We demonstrate that TAL is less suited for the immediate classification of small-sized windows of data, yet offers an interesting perspective on inertial-based HAR – alleviating the need for fixed-size windows and enabling algorithms to recognize activities of arbitrary length. With design choices and training concepts yet to be explored, we argue that TAL architectures could be of significant value to the inertial-based HAR community. The code and data download to reproduce experiments is publicly available via github.com/mariusbock/tal_for_har.
@@ 1. Introduction
The recognition of performed activities through wearable sensors such as Inertial Measurement Units (IMUs) has shown to be of significant value in areas such as health care or the support of complex work processes (Bulling et al., 2014). With the works of many researchers exploring lightweight architectures (Zhou et al., 2022), much of the success of inertial-based models has stemmed from applying them in an online-fashion on wearable edge devices (Ravi et al., 2005; Ward et al., 2006). As shown to work on a multitude of HAR-related scenarios, it has established itself within the inertial-based community to employ a window-based prediction approach using a predefined window size and overlap. With classification algorithms being tasked to assign a label to each window individually, this approach has enabled inertial-based architectures to classify newly incoming data at will. As a too large window size may include multiple fast actions within a single window, and a too short window may not be enough to capture a complete lengthier action, the chosen size of the fixed window has become a crucial parameter within traditional recognition systems.
@@
A persistent trend in Deep Learning has been the applicability of machine learning concepts such as self-attention (Vaswani et al., 2017) to other areas and application scenarios than originally introduced for. With significant progress having been made since the introduction of deep neural architectures such as DeepConvLSTM (Ordóñez and Roggen, 2016), researchers have followed this trend and have continuously worked on improving the architectural design of networks by incorporating newly introduced techniques (see, e.g., (Zhou et al., 2022)). A promising recent approach in video-based Human Activity Recognition (HAR) is Temporal Action Localization (TAL), which aims to locate activity segments, defined by a class label, start, and end point, within an untrimmed video. Even though introduced architectures have almost doubled in performance over the last 5 years on existing datasets like THUMOS-14 (Jiang et al., 2014), results on large corpora such as EPIC-KITCHENS-100 (Damen et al., 2022) and Ego4D (Grauman et al., 2022) show that the prediction problem is far from being saturated.
Despite sharing a mutual goal, the TAL community as opposed to the inertial community does not follow an online-based prediction approach, but aims to classify recorded times sequences in an offline fashion, taking advantage of learnable temporal dependencies along the complete timeline. Furthermore, by following a segment-based prediction with activity boundaries being learned via regression-based loss, the ”window size problem” is alleviated with methods being flexible enough to recognize both short and long lasting activities. TAL models have recently been shown to be capable of being trained using raw inertial data (Bock et al., 2023), marking the first instance of such vision-based models being applied in the context of inertial-based HAR. When compared with popular inertial models, results on one particular dataset have already shown that TAL models can produce more coherent and less fragmented predictions while maintaining performance in terms of traditional classification metrics. In light of activity recognition systems having also been deployed to provide offline analysis of streams of prerecorded data (Inoue et al., 2019; Van Laerhoven et al., 2008; Debes et al., 2016; Blanke and Schiele, 2009; Zhai et al., 2021; Berchtold et al., 2010), this work sets out to further examine the applicability of Temporal Action Localization for inertial-based HAR. Our contributions in this paper are three-fold:
@@
    We demonstrate the capabilities of a novel approach to inertial-based HAR, being single-stage TAL, to outperform popular, window-based inertial models on a multitude of wearable activity recognition benchmark datasets in an offline-prediction scenario.
@@
    Complementing traditional, window-wise classification metrics, we introduce a set of unexplored, segment-based evaluation metrics for inertial-based HAR, which are based around the scalar evaluation metric mean Average Precision (Alwassel et al., 2018).
@@
    Though our results demonstrate the superiority of inertial-based models in an online prediction scenario, we show that TAL models can be applied in a near-online fashion, functioning e.g. a server-side prediction tool which provides prediction with a lag of 1 minute.
@@ 2. Related Work
With on-body sensors providing a robust and non-intrusive way to monitor participants along long stretches of time, research conducted in the area of inertial-based HAR has worked towards the automatic interpretation of one or multiple sensor streams to reliably detect activities e.g. in the context of providing medical support or providing guidance during complex work processes (Bulling et al., 2014). With deep neural networks having established themselves as the de facto standard in inertial-based HAR, DeepConvLSTM (Ordóñez and Roggen, 2016) as well as the models introduced at a later point follow a similar prediction scenario design applying a sliding window approach which groups concurrent data points for classification. In their original publication Ordóñez and Roggen (2016) have found a combination of convolutional and recurrent layers to produce competitive results, with the idea being to model temporal dependencies amongst automatically extracted discriminative features within a sliding window in order to classify it correctly as one of the $N$ activity classes or, if applicable, NULL-class. Building on the idea of combining these two types of layers, researchers have worked on extending the original DeepConvLSTM or have introduced their own architectural designs (Murahari and Plötz, 2018; Vaizman et al., 2018; Peng et al., 2018; Xi et al., 2018; Xu et al., 2019; Xia et al., 2020; Liu et al., 2020; Chen et al., 2020; Bock et al., 2021; Abedin et al., 2021; Zhou et al., 2022; Zhang et al., 2022a; Miao et al., 2022; Chen et al., 2023b, a; Wieland and Pankratius, 2023). A simple modification of the DeepConvLSTM is the shallow DeepConvLSTM (Bock et al., 2021). Contradicting the popular belief that one needs at least two recurrent layers when dealing with time series data (Karpathy et al., 2015), Bock et al. (2021) demonstrated that removing the second LSTM layer within the original DeepConvLSTM architecture results in significant increases in predictive performance on a multitude of HAR benchmark datasets while also decreasing the number of learnable parameters. Furthermore, with the original DeepConvLSTM only being able to learn temporal dependencies within a sliding window, the shallow DeepConvLSTM applies the LSTM across batches, effectively making the batches the sequence which is to be learned by the LSTM. This dimension flip, along with a non-shuffled training dataset, enables the architecture to learn temporal dependencies amongst a batched input. The same year as the shallow DeepConvLSTM, Abedin et al. (2021) introduced Attend-and-Discriminate, a deep neural network architecture following the idea of the original DeepConvLSTM by combining both convolutional and recurrent layers, yet further adding a cross-channel interaction encoder using self-attention as well as attention mechanism to the recurrent parts of the network. In 2022 Zhou et al. (2022) proposed TinyHAR, a lightweight HAR model that uses a transformer encoder block along with means of cross-channel fusion via a fully connected layer and a final self-attention layer which aims to learn the contribution of each outputted time step produced by the recurrent parts individually.
@@
Classifying videos in the context of Human Activity Recognition can be broadly categorized into three main application scenarios: Action Recognition, which aims to classify trimmed videos into one of $C$ activity classes; Action Anticipation, which aims to predict the next likely activities after observing a preceding video sequence; and Temporal Action Localization (TAL), which seeks to identify and locate all activity segments within an untrimmed video. With the inertial-based benchmark datasets consisting of a multitude of activities, TAL is to be considered most comparable to inertial-based HAR. Unlike popular intertial architectures though, TAL models aim to predict all segments within an untrimmed video at once. Existing TAL methods can broadly be categorized into two categories: two-stage and one-stage recognition systems. Two-stage recognition system (Lin et al., 2019, 2020; Xu et al., 2020; Bai et al., 2020; Zhao et al., 2020; Zeng et al., 2019; Gong et al., 2020; Liu et al., 2021; Qing et al., 2021; Sridhar et al., 2021; Zhu et al., 2021; Zhao et al., 2021; Tan et al., 2021) divide the process of identifying actions within an untrimmed video into two subtasks. First, during proposal generation, candidate segments are generated, which are then, during the second step, classified into one of $C$ activity classes and iteratively refined regarding their start and end times. Contrarily, single-stage models (Yang et al., 2023; Shi et al., 2022; Nag et al., 2022; Liu et al., 2022; Liu and Wang, 2020; Long et al., 2019; Lin et al., 2021; Chen et al., 2022; Zhang et al., 2022b; Tang et al., 2023; Shi et al., 2023) do not rely on activity proposals, directly aiming to localize segments along their class label and refined boundaries. In 2022 Zhang et al. (2022b) introduced the ActionFormer, a lightweight, single-stage TAL model which unlike previously introduced single-stage architectures does not rely on pre-defined anchor windows. In line with the success of transformers in other research fields, Zhang et al. (2022b) demonstrated their applicability for TAL, outperforming previously held benchmarks on several TAL datasets (Heilbron et al., 2015; Damen et al., 2022; Jiang et al., 2014) by a significant margin. Surprisingly, a year later, TemporalMaxer suggested removing transformer layers within the ActionFormer, arguing that feature embeddings are already discriminative enough (Tang et al., 2023). Though being more lightweight than the ActionFormer, the TemporalMaxer showed to outperform its precedent during benchmark analysis. Similarly to the TemporalMaxer, Shi et al. (2023) introduced TriDet, which suggested altering ActionFormer in two ways. First, to mitigate the risk of rank-loss, self-attention layers are replaced with SGP layers. Second, the regression head in the decoder is replaced with a Trident head, which improves imprecise boundary predictions via an estimated relative probability distribution around each timestamp’s activity boundaries.
@@ 3. Temporal Action Localization for Inertial-based HAR
As the inertial-based HAR and TAL communities deal with inherently different modalities, both communities have developed distinct predictive pipelines and algorithms tailored to the challenges and characteristics of their respective modalities (see Figure 1). The objective of both inertial activity recognition and TAL is to predict all activities present in an untrimmed timeline. Given an input data stream $X$ of a sample participant, both the inertial and TAL communities start by applying a sliding window approach which shifts over $X$, dividing the input data into windows, e.g. of one second duration with a 50% overlap between consecutive windows. This process results in $X=\{\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{T}\}$ being discretized into $t=\{0,1,...,T\}$ time steps, where $T$ is the number of windows for each participant. It is important to note that the TAL models do not use raw data as input but are instead trained using feature embeddings extracted from each individual sliding video clip, which are extracted using pre-trained methods such as (Carreira and Zisserman, 2017).
Given all sliding windows associated with an untrimmed sequence, inertial activity recognition models aim to predict an activity label $a_{t}$ for each sliding window $\mathbf{x}_{t}$, where $a$ belongs to a predefined set of activity labels, $a_{t}\in{1,...,C}$. To do so, the sliding windows are batched and fed through, e.g., a deep neural network, such as the DeepConvLSTM (Ordóñez and Roggen, 2016). The resulting activity labels for each window are then compared to the true activity labels from the ground truth data, and classification metrics like accuracy or F1-score are calculated. Contrarily, TAL models aim to identify and localize segments of actions within the untrimmed data stream, which can span across multiple windows. To achieve this algorithms are trained to predict activity segments $Y=\{\mathbf{y}_{1},\mathbf{y}_{2},...,\mathbf{y}_{N}\}$, where $N$ varies across participants. Each activity instance $\mathbf{y_{i}}=(s_{i},e_{i},a_{i})$ is defined by its starting time $s_{i}$ (onset), end time $e_{i}$ (offset) and its associated activity label $a_{i}$, where $s_{i}\in[0,T]$, $e_{i}\in[0,T]$, and $a_{i}\in\{1,...,C\}$.
As the input data used to train TAL models is a collection of 1-dimensional feature embeddings, the 2 d-dimensional, windowed inertial data commonly found in the inertial-based HAR community, cannot directly be used to train TAL models. The following will thus describe two preprocessing methods which can be used to convert the inertial data into a format such that it can be used as input for TAL models.
Vectorization of raw inertial data
Since both communities employ a sliding window approach but feed data to their models using different dimensions, Bock et al. (2023) proposed a simple, yet effective preprocessing step. This step converts the 2-dimensional, windowed inertial data, as used by inertial architectures, into a 1-dimensional feature embedding suitable for training TAL models. The preprocessing method as detailed in Figure 2 involves concatenating the different sensor axes of each window, converting the input data to be a collection of 1-dimensional feature embedding vectors $\mathbf{x}_{t}\in\mathbb{R}^{1\times WS}$ where $W$ is the number of samples within a window, and $S$ is the number of sensor axes. More specifically, given a sliding window $\mathbf{x}_{SW}\in\mathbb{R}^{W\times S}$, we vectorize the two-dimensional matrix as follows:
@@
[…]
A visualization of the prediction pipeline of inertial-based activity recognition and single-stage Temporal Action Localization. Within inertial-based activity recognition inertial data, illustrated as three sensor streams, is chunked into sliding windows which are then predicted as one of the activity classes by a Deep Learning model. The predicted labels are then compared to that of the ground truth and classification metrics are calculated. Temporal Action Localization first extracts video embeddings from sliding video clips. The embeddings are then fed through a Deep Learning model which predicts activity segments, visualized as colored boxes on a timeline, that are in a final step compared for overlap and label correctness with the ground truth.
@@
[…]
Visualization of the vecotrization of a sliding window of the inertial data such that it can be used as input for TAL models. For each window each of the three axis of a sensors are grouped together (i.e., a 2-dimensional matrix). The vectorization applied in this paper concatenates each axis of each sensors one after another such that the each two-dimensional sliding window is converted to a 1-dimensional feature embedding. Within the example 12 axis with 50 samples each are thus converted to a feature vector of size 600.
Two-stage training via prepended inertial models
In order to encode videos and come up with a discriminative, latent feature representation, TAL models usually resort to using extracted feature embeddings from models pretrained on large vision corpuses like Kinetics-400 (Kay et al., 2017). Inspired by this, we propose a second variant on how to use inertial data as input to TAL models (see Figure 3) which involves using features extracted from a separately trained inertial model as latent representation of each sliding window. Specifically, a LOSO cross-validation step within the two stage training consists of 1. training the inertial model on all but the validation data, 2. use the trained inertial network to extract latent feature representations of each window within the training and validation data, and 3. using the extracted features as clip-wise feature embeddings to train a TAL model.
Figure 3. Visualization of the applied two-stage training process. The first stage involves training e.g. a classic DeepConvLSTM as introduced by Ordóñez and Roggen (2016). Once the first-stage training has finished, the classifier is omitted from the model such that latent features can be extracted. The 1-dimensional, window-wise features are then used as input embeddings for the second stage, i.e. training a TAL model.
[…]
Visualization of the applied two-stage training process. The first stage involves training e.g. a classic DeepConvLSTM as introduced by Ordóñez and Roggen (2016). Once the first-stage training has finished, the classifier is omitted from the model such that latent features can be extracted. The 1-dimensional, window-wise features are then used as input embeddings for the second stage, i.e. training a TAL model.
3.1. TAL Architectures Overview
In light of the recent success of transformer-based models in e.g. Natural Language Processing and Computer Vision, Zhang et al. (2022b) proposed the ActionFormer, is an end-to-end trainable transformer-based TAL architecture. Unlike other single-stage TAL approaches, it does not rely on pre-defined anchor windows. The architecture, as illustrated in Figure 4, combines multiscale feature representations with local self-attention and is trained through a classification and regression loss calculated by a light-weighted decoder. Building up on the ActionFormer architecture, Tang et al. (2023) and Shi et al. (2023) proposed the TemporalMaxer and TriDet model respectively. Within the TriDet model projection and transformer layers of the ActionFormer are replaced with fully-convolutional SGP layers and the regression head is replaced by a trident regression head which claims to improve imprecise boundary predictions. Contrarily, the TemporalMaxer suggests modifying the encoder of the ActionFormer to employ solely max pooling and remove all transformer-based layers, as, according to the authors, this does not come at the cost of a lost in information and predictive performance.